For all even integral values of n , lim
x-infinite (X+1)^-n exists
1
Expert's answer
2020-03-13T12:10:37-0400

In general the limit does not exist.

Case:1 If "n>0" .

In this case, "\lim_{x\to \infty}(x+1)^{-n}=\lim_{x\to \infty} \frac{1}{(x+1)^n}" "=0"

Hence the limit exists.

Case: 2 If "n=0"

In this case, "\lim_{x\to \infty}{(x+1)^{-n}}=1"

Hence the limit exists.

Case:3 If "n<0"

In this case, "\lim_{x\to \infty}(x+1)^{-n}=\infty"

And the limit does not exist.